# Breaking the Gaussian Barrier: Residual-PAC Privacy for Automatic Privatization

### v1

@@ 2.1 Automatic PAC Privatization Algorithms
This result establishes a simple upper bound on the mutual information after perturbation and implies that the noise can be tailored anisotropically. To apply this bound in practice, the covariance $\Sigma_{\mathcal{M}(X)}$ must be estimated from Monte Carlo simulations. The authors of [1] propose a high-confidence noise calibration protocol (Algorithm 1) that determines an appropriate noise covariance $\Sigma_{B}$ to ensure that $\mathtt{MI}(X;\mathcal{M}(X)+B)\leq v+\beta$ with confidence at least $1-\gamma$, given user-specified parameters $v$, $\beta$, and sample complexity $m$. We refer to $\widehat{\beta}=v+\beta$ as the privacy budget. The framework extends naturally to randomized mechanisms of the form $\mathcal{M}(X,\theta)$, where $\theta$ is a random seed (Corollary 2 of [1]).
Recent work by Sridhar et al. [15] improves the practicality and efficiency of PAC Privacy (i.e., Efficient-PAC) by introducing an anisotropic noise calibration scheme that avoids full covariance estimation (Algorithm 2). Instead of computing the entire output covariance matrix and performing expensive matrix decomposition, their method projects mechanism outputs onto a unitary basis and estimates only the per-direction variances. This leads to a more scalable and sample-efficient algorithm while still ensuring rigorous mutual information guarantees under the PAC Privacy framework.
In addition, the authors address a key limitation of black-box privacy mechanisms—namely, the instability of outputs caused by random seeds, arbitrary encodings, or non-deterministic implementations. They propose methods for reducing such instability through output regularization and canonicalization, enabling more consistent noise calibration and better overall utility. These refinements are particularly impactful in high-dimensional or structure-sensitive learning tasks, where the original PAC scheme may incur unnecessary noise due to variability not intrinsic to the learning objective.
Theorem 2 (Theorem 1 of [15]).
Algorithm 2 Efficient-PAC [15]
2:  while $m\leq 2$ or $f_{\tau}(\bm{\sigma}_{m-1},\bm{\sigma}_{m})\geq\tau$ do
3:     Sample $X_{m}\sim\mathcal{D}$, compute $y_{m}\leftarrow\mathcal{M}(X_{m})$
6:  end while
7:  for $i=1$ to $d$ do
9:  end for
10:  return  $\Sigma_{B}$ with $\Sigma_{B}[i][i]=e_{i}$
@@ 3 Residual-PAC Privacy
PAC Advantage Privacy, as defined in Definition 3, quantifies the amount of privacy leakage—denoted by $\Delta_{f}^{\delta}$—incurred from releasing $\mathcal{M}(X)$, measured as the (prior-expected)posterior advantage, using $f$-divergence, gained by the adversary after observing $\mathcal{M}(X)$. Complementing this perspective, we introduce the notion of posterior disadvantage encountered by the adversary, which captures the amount of residual privacy protection that persists after information has been leaked through the mechanism’s output. To formalize this residual protection, we first define the intrinsic privacy of a data distribution $\mathcal{D}$ as its closeness to the uniform distribution over $\mathcal{X}$. Specifically, let $\mathcal{U}$ denote the uniform distribution on $\mathcal{X}$, and define
$\mathtt{IntP}_{f}(\mathcal{D})=-\,\mathtt{D}_{f}(\mathcal{D}\,\|\,\mathcal{U}),$
@@
where $\mathtt{D}_{f}(\mathcal{D}\|\mathcal{U})$ is the $f$-divergence between $\mathcal{D}$ and $\mathcal{U}$, quantifying how much $\mathcal{D}$ deviates from perfect uniformity. Intuitively, $-\mathtt{D}_{f}(\mathcal{D}\|\mathcal{U})$ serves as a reward for uniformity (i.e., maximal intrinsic privacy), and by construction, $\mathtt{IntP}_{f}(\mathcal{D})\leq 0$, attaining zero only when $\mathcal{D}$ is uniform.
Definition 5 ($(\mathtt{R}_{f}^{\delta},\rho,\mathcal{D})$ Residual-PAC Privacy).
A mechanism $\mathcal{M}$ is said to be $(\mathtt{R}_{f}^{\delta},\rho,\mathcal{D})$ Residual-PAC private if it is $(\delta,\rho,\mathcal{D})$ PAC private and
$\mathtt{R}_{f}^{\delta}\equiv\mathtt{IntP}_{f}(\mathcal{D})-\mathtt{D}_{f}(\mathbf{1}_{\delta}\|\mathbf{1}_{\delta_{\rho}^{o}}),$
@@
is the posterior disadvantage, where $\mathbf{1}_{\delta}$ and $\mathbf{1}_{\delta_{\rho}^{o}}$ are indicator distributions representing the adversary’s inference success before and after observing the mechanism’s output, respectively.
@@
Under this definition, $\mathtt{R}_{f}^{\delta}$ characterizes the residual privacy guarantee—that is, the portion of intrinsic privacy that remains uncompromised after the adversary’s inference. Notably, the total intrinsic privacy is precisely decomposed as
$\mathtt{IntP}_{f}(\mathcal{D})=\mathtt{R}_{f}^{\delta}+\Delta_{f}^{\delta},$
@@
where $\Delta_{f}^{\delta}$ is the PAC Advantage Privacy, or privacy leakage, term. This relationship provides a complete and interpretable accounting of privacy risk, distinguishing between the privacy that is lost and that which endures after information disclosure.

### v2

@@ 2.2 Automatic PAC Privatization Algorithms
Theorem 1 establishes a simple upper bound on the mutual information with Gaussian noise perturbation. Choosing $\Sigma_{B}$ to implement the Gaussian surrogate bound $\mathtt{LogDet}(\mathcal{M}(X),B)=\beta$ for a privacy budget $\beta$ enables anisotropic noise as it estimates the eigenvectors of $\mathcal{M}(X)$ to fit the instance-based noise to the geometry of the eigenspace of $\mathcal{M}(X)$. The result extends naturally to randomized mechanisms of the form $\mathcal{M}(X,\theta)$, where $\theta$ is a random seed (Corollary 2 of [46]). Building on Theorem 1, an automatic PAC privatization algorithm (Auto-PAC) shown in Algorithm 1 is proposed by [46] to determine an appropriate Gaussian noise covariance $\Sigma_{B}$ to ensure that $\mathtt{MI}(X;\mathcal{M}(X)+B)\leq\beta$ with confidence at least $1-\gamma$. This is achieved using the user-specific security parameter $c$, privacy budget partitions $v$ and $\beta^{\prime}$ such that $\beta=v+\beta^{\prime}$ (Theorem 4 of [46]). We refer to Algorithm 1 as $(1-\gamma)$-Confidence Auto-PAC.
2.3 Differential Privacy
In addition to the standard PAC Privacy, we also compare our approach to the differential privacy (DP) framework. Let $X$ be the input dataset. Each data point $x_{i}$ is defined over some measurable domain $\mathcal{X}^{\dagger}$, so that $x=(x_{1},x_{2},\ldots,x_{n})\in\mathcal{X}=(\mathcal{X}^{\dagger})^{n}$. We say two datasets $x,x^{\prime}\in\mathcal{X}$ are adjacent if they differ in exactly one data point.
Definition 5 ($(\epsilon,\bar{\delta})$-Differential Privacy [14]).
A randomized mechanism $\mathcal{M}:\mathcal{X}\mapsto\mathcal{Y}$ is said to be $(\epsilon,\bar{\delta})$-differentially private (DP), with $\epsilon\geq 0$ and $\bar{\delta}\in[0,1]$, if for any pair of adjacent datasets $x,x^{\prime}$, and any measurable $\mathcal{W}\subseteq\mathcal{Y}$, it holds that $\Pr[\mathcal{M}(x)\in\mathcal{W}]\leq e^{\epsilon}\Pr[\mathcal{M}(x^{\prime})\in\mathcal{W}]+\bar{\delta}$.
The parameter $\epsilon$ is usually referred to as the privacy budget, and $\delta\in(0,1]$ represents the failure probability. DP is an input-independent adversarial worst-case approaches that focus on the sensitivity magnitude, while Auto-PAC is instance-based and adds anisotropic noise tailored to each direction as needed. Appendix C characterizes the difference between DP, PAC Privacy, and our Residual-PAC (R-PAC) Privacy.
This section characterizes the utility of Auto-PAC by focusing on the conservativeness of the implemented mutual information bounds. To distinguish from Algorithm 1 ($(1-\gamma)$-confidence Auto-PAC), we use Auto-PAC (algorithm) to refer to the direct implementation of privacy budgets for the bound $\mathtt{LogDet}(\mathcal{M}(X),B)$ without a target conference level.
The Gaussian surrogate bound is conservative due to a nonzero Gaussianity gap, the discrepancy between the true mutual information and $\mathtt{LogDet}(\mathcal{M}(X),B)$ defined by (2):
$\mathtt{Gap}_{\mathtt{d}}\equiv\mathtt{LogDet}(\mathcal{M}(X),B)-\mathtt{MI}(X;\mathcal{M}(X)+B).$ (3)
Define $Z=\mathcal{M}(X)+B$ with mean $\mu_{Z}=\mu_{\mathcal{M}(X)}$ and covariance $\Sigma_{Z}=\Sigma_{\mathcal{M}(X)}+\Sigma_{B}$. Let $P_{\mathcal{M},B}$ denote the true distribution of $Z=\mathcal{M}(X)+B$, and define the Gaussian surrogate distribution as
$\widetilde{Q}_{\mathcal{M}}\equiv\mathcal{N}(\mu_{Z},\Sigma_{Z})$ (4)
with the same first and second moments as $Z\sim P_{\mathcal{M},B}$.
Proposition 1.
Let $B\sim\mathcal{N}(0,\Sigma_{B})$. Then, $\mathtt{Gap}_{\mathtt{d}}=\mathtt{D}_{\mathrm{KL}}(P_{\mathcal{M},B}\|\widetilde{Q}_{\mathcal{M}})\geq 0$. Moreover, $\mathtt{Gap}_{\mathtt{d}}=0$ iff $P_{\mathcal{M},B}=\widetilde{Q}_{\mathcal{M}}$.
[…]
Proposition 2.
For any privacy budget $\beta>0$, the noise distribution $Q=\mathcal{N}(0,\Sigma_{B})$ obtained by Auto-PAC is the unique solution of the following problem:
$\inf_{B\sim Q^{\prime}}\mathbb{E}_{Q^{\prime}}\bigl{[}\|B\|_{2}^{2}\bigr{]}\quad\text{s.t.}\quad\mathtt{MI}(X;\widetilde{Z})\leq\beta\text{ with }\widetilde{Z}\sim\widetilde{Q}_{\mathcal{M}}$. (5)
Proposition 2 implies that Auto-PAC’s zero-mean Gaussian noise is the optimal solution to minimize the magnitude of the noise subject to the mutual information constraint if we replace $Z\sim P_{\mathcal{M},B}$ by $\widetilde{Z}\sim\widetilde{Q}_{\mathcal{M}}$.
Proposition 3.
For any privacy budget $\beta>0$, let $Q$ and $Q_{\gamma}$, respectively, be the Gaussian noise distribution obtained by Auto-PAC and $(1-\gamma)$-Confidence Auto-PAC with any $\gamma\in[0,1]$. Let $B\sim Q$ and $B_{\gamma}\sim Q_{\gamma}$. Then, the following holds.
  • (i)
    $\mathtt{MI}(X;\mathcal{M}(X)+B_{\gamma})\leq\mathtt{MI}(X;\mathcal{M}(X)+B)$.
  • (ii)
    $\mathbb{E}_{Q_{\gamma}}[\|B_{\gamma}\|^{2}_{2}]\geq\mathbb{E}_{Q}[\|B\|^{2}_{2}].$
In Proposition 3, part (i) shows that $(1-\gamma)$-confidence Auto-PAC is more conservative than directly implementing $\mathtt{LogDet}(\mathcal{M}(X),B)$ (Auto-PAC) for the same privacy budget. Part (ii) demonstrates that $(1-\gamma)$-confidence Auto-PAC uses larger noise magnitude than Auto-PAC for the same privacy budget. Thus, in subsequent comparisons involving PAC Privacy, we focus on Auto-PAC.
3.1 Mechanism Comparison in PAC Privacy
Definition 9 of [46] defines the optimal perturbation that tightly implements the privacy budget while maintaining optimal utility, where utility is captured by a loss function $\mathcal{K}$. An optimal perturbation $Q^{*}$ is a solution of the following optimization problem:
$\inf_{Q}\mathbb{E}_{Q,\mathcal{M},\mathcal{D}}[\mathcal{K}(B;\mathcal{M})]\quad\text{s.t.}\quad\mathtt{MI}(X;\mathcal{M}(X)+B)\leq\beta,\;B\sim Q.$ (6)
The choice of utility loss function $\mathcal{K}$ is context-dependent. However, in many applications, we are primarily concerned with the expected Euclidean norm of the noise or a convex function thereof, e.g., $\mathbb{E}_{Q,\mathcal{M},\mathcal{D}}[\mathcal{K}(B;\mathcal{M})]=\mathbb{E}_{Q}\bigl{[}\|B\|_{2}^{2}\bigr{]}$.
We now show that using $\mathbb{E}_{Q,\mathcal{M},\mathcal{D}}[\mathcal{K}(B;\mathcal{M})]=\mathbb{E}_{Q}\bigl{[}\|B\|_{2}^{2}\bigr{]}$ is sufficient to obtain perturbations that maintain coherent ordering of PAC Privacy using mutual information (i.e., larger privacy budgets yield non-decreasing actual mutual information).
Proposition 4.
Fix a mechanism $\mathcal{M}$ and data distribution $\mathcal{D}$. Let $\mathcal{Q}$ denote the collection of all zero-mean noise distributions under consideration, and let $\mathtt{I}_{\mathrm{true}}:\mathcal{Q}\mapsto\mathbb{R}_{\geq 0}$ be the true mutual information functional; i.e., $\mathtt{I}_{\mathrm{true}}(Q)=\mathtt{MI}(X;\mathcal{M}(X)+B)$ with $B\sim Q$ for $Q\in\mathcal{Q}$. For each privacy budget $\beta\geq 0$, define the feasible region $\mathcal{F}(\beta)\equiv\{Q\in\mathcal{Q}:\mathtt{I}_{\mathrm{true}}(Q)\leq\beta\}.$ Suppose that $\mathcal{F}(\beta)$ is nonempty for all privacy budgets of interest. For each $\beta\geq 0$, let $Q^{*}(\beta)$ be a solution of the problem:
$\min_{B\sim Q}\mathbb{E}_{Q}[\|B\|^{2}_{2}]\quad\text{s.t.}\quad Q\in\mathcal{F}(\beta).$ (7)
However, if Auto-PAC is used to solve the optimization problem (5), we have the conservative implementation of a given privacy budget. The next result shows that when $\mathtt{Gap}_{\mathtt{d}}=\mathtt{D}_{\mathrm{KL}}(P_{\mathcal{M},B}\|\widetilde{Q}_{\mathcal{M}})>0$, Auto-PAC does not, in general, maintain coherent ordering of PAC Privacy.
[…]
Theorem 2.
Fix a mechanism $\mathcal{M}$ and data distribution $\mathcal{D}$. Let $\mathcal{Q}$ denote the collection of all zero-mean noise distributions under consideration, and let $\mathtt{I}_{\mathrm{true}}:\mathcal{Q}\mapsto\mathbb{R}_{\geq 0}$ be the true mutual information functional; i.e., $\mathtt{I}_{\mathrm{true}}(Q)=\mathtt{MI}(X;\mathcal{M}(X)+B)$ with $B\sim Q$ for $Q\in\mathcal{Q}$. For each $\beta\geq 0$, let $Q^{*}(\beta)$ be a solution of the optimization in Proposition 2. For any $0<\beta_{1}<\beta_{2}$, define $\mathtt{G}(\beta_{2},\beta_{1})\equiv\mathtt{Gap}_{\mathtt{d}}(Q^{*}(\beta_{2}))-\mathtt{Gap}_{\mathtt{d}}(Q^{*}(\beta_{1})).$ Then:
  1. (i)
  2. (ii)
Theorem 2 characterizes when Auto-PAC maintains coherent ordering of actual information leakage $\mathtt{I}_{\mathrm{true}}=\beta-\mathtt{Gap}_{\mathtt{d}}$. Increasing the budget from $\beta_{1}$ to $\beta_{2}$ permits extra leakage $\beta_{2}-\beta_{1}$ by using Auto-PAC, but part may be wasted if the mechanism output becomes more non-Gaussian. The wasted portion is $\mathtt{G}(\beta_{2},\beta_{1})=\mathtt{Gap}_{\mathtt{d}}(Q^{*}(\beta_{2}))-\mathtt{Gap}_{\mathtt{d}}(Q^{*}(\beta_{1}))$. If this waste exceeds the budget increase, then $\mathtt{I}_{\mathrm{true}}$ decreases despite a larger nominal budget, violating coherent ordering. This result cautions against comparing mechanisms using Auto-PAC solely by budgets, as identical budgets may yield different true leakages depending on their respective Gaussianity gaps.
3.2 $\mathtt{Gap}_{\mathtt{d}}$ Reduction via Non-Gaussianity Correction
[…]
For any estimator $\widehat{\mathtt{D}}_{Z}$ of $\mathtt{D}_{Z}$, define the improved mutual information estimate:
For $0leq\widehat{\mathtt{D}}_{Z}\leq\mathtt{D}_{Z}$, we have
$\mathtt{MI}(X;\mathcal{M}(X)+B)\leq\mathtt{IMI}(\widehat{\mathtt{D}}_{Z})\leq\mathtt{LogDet}(\mathcal{M}(X),B).$
Thus, if we can obtain $\widehat{\mathtt{D}}_{Z}$ satisfying $0\leq\widehat{\mathtt{D}}_{Z}\leq\mathtt{D}_{Z}$ after Auto-PAC privatization, then for any $\Sigma_{B}$ that ensures $\mathtt{LogDet}(\mathcal{M}(X),B)=\beta$, we have $\mathtt{IMI}(\widehat{\mathtt{D}}_{Z})=\beta-\widehat{\mathtt{D}}_{Z}$ as surrogate upper bound that is tighter than $\mathtt{LogDet}(\mathcal{M}(X),B)$. Thus, we can have tighter privacy accounting post-hoc to the Auto-PAC privatization to save additional privacy budget, without requiring direct mutual information estimation.
Before describing the approaches, we first introduce two standard discrepancy measures between $P_{\mathcal{M},B}$ and $\widetilde{Q}_{\mathcal{M}}$.
Definition 6 (Donsker–Varadhan (DV) Objective [11]).
For probability measures $P$ and $Q$ on a common measurable space,
$\mathtt{D}_{\mathrm{KL}}(P\|Q)\;=\;\sup_{f:\,\mathcal{Y}\to\mathbb{R}}\Big{\{}\,\mathbb{E}_{P}[f(Y)]-\log\mathbb{E}_{Q}\big{[}e^{f(Y)}\big{]}\Big{\}},$
where the supremum ranges over measurable $f$ such that $\mathbb{E}_{Q}[e^{f}]<\infty$. We call $\mathcal{J}(f;P,Q)\equiv\mathbb{E}_{P}[f]-\log\mathbb{E}_{Q}[e^{f}]$ the DV objective. In our setting, $\mathtt{D}_{Z}=\mathtt{D}_{\mathrm{KL}}(P_{Z}\|\widetilde{Q}_{\mathcal{M}})=\sup_{f}\mathcal{J}\big{(}f;P_{\mathcal{M},B},\widetilde{Q}_{\mathcal{M}}\big{)}$.
Definition 7 (Sliced Wasserstein Distances (SWD)).
For $p\geq 1$, the $p$-Wasserstein distance between $P$ and $Q$ on $\mathbb{R}^{d}$ is
$W_{p}(P,Q)\;=\;\Big{(}\inf_{\eta\in\widehat{\Pi}(P,Q)}\,\mathbb{E}_{(X,Y)\sim\eta}\big{[}\|X-Y\|_{2}^{p}\big{]}\Big{)}^{1/p},$
where $\widehat{\Pi}(P,Q)$ is the set of couplings with marginals $P$ and $Q$. The sliced $p$-Wasserstein distance averages 1-D Wasserstein distances over directions $v$ on the unit sphere $\mathbb{S}^{d-1}$:
$\mathrm{SW}_{p}^{p}(P,Q)\;=\;\int_{\mathbb{S}^{d-1}}W_{p}^{p}\big{(}\mathcal{L}(\langle v,X\rangle),\,\mathcal{L}(\langle v,Y\rangle)\big{)}\,d\sigma(v),$
[…]
In addition, we define the following finite-sample lower-confidence estimator.
Definition 8 (Finite-Sample Lower-Confidence DV Estimator).
Fix a function class $\mathcal{F}\subset\{f:\mathbb{R}^{d}\to\mathbb{R}\}$ with $0\in\mathcal{F}$ and let $\widehat{\mathcal{J}}(f;S_{P},S_{Q})\;\equiv\;\frac{1}{|S_{P}|}\sum_{z\in S_{P}}f(z)\;-\;\log\Big{(}\frac{1}{|S_{Q}|}\sum_{z\in S_{Q}}e^{f(z)}\Big{)}$ denote the empirical DV objective on samples $S_{P}$ from $P_{Z}$ and $S_{Q}$ from $\widetilde{Q}_{\mathcal{M}}$. Draw four independent splits $S_{P}^{\mathrm{tr}},S_{Q}^{\mathrm{tr}},S_{P}^{\mathrm{val}},S_{Q}^{\mathrm{val}}$ with sizes $n_{P}^{\mathrm{tr}},n_{Q}^{\mathrm{tr}},n_{P}^{\mathrm{val}},n_{Q}^{\mathrm{val}}$ respectively, and fit $\widehat{f}_{\mathrm{tr}}\;\in\;\arg\max_{f\in\mathcal{F}}\;\widehat{\mathcal{J}}\big{(}f;S_{P}^{\mathrm{tr}},S_{Q}^{\mathrm{tr}}\big{)}.$
[…]
$\widehat{\mathtt{D}}_{\mathrm{LCE}}\equiv\Big{[}\widehat{\mathcal{J}}\big{(}\widehat{f}_{\mathrm{tr}};S_{P}^{\mathrm{val}},S_{Q}^{\mathrm{val}}\big{)}-\Gamma_{\hat{\delta}}\Big{]}_{+}$.
Theorem 3 (DV-Based Correction).
Let $Z=\mathcal{M}(X)+B$ with deterministic $\mathcal{M}$ and $B\sim\mathcal{N}(0,\Sigma_{B})$, and let $\widetilde{Q}_{\mathcal{M}}=\mathcal{N}(\mu_{Z},\Sigma_{Z})$ be the Gaussian surrogate defined in (4). Assume $P_{\mathcal{M},B}\ll\widetilde{Q}_{\mathcal{M}}$. For any measurable $f:\mathbb{R}^{d}\to\mathbb{R}$ with $\mathbb{E}_{\widetilde{Q}_{\mathcal{M}}}[e^{f(Z)}]<\infty$, define
$\widehat{\mathtt{D}}_{Z}(f)\equiv\mathcal{J}\big{(}f;P_{\mathcal{M},B},\widetilde{Q}_{\mathcal{M}}\big{)}=\mathbb{E}_{P_{Z}}[f(Z)]-\log\mathbb{E}_{\widetilde{Q}_{\mathcal{M}}}\big{[}e^{f(Z)}\big{]}.$
Let $\widehat{\mathtt{D}}_{\mathrm{LCE}}$ be the finite-sample lower-confidence estimator from Definition 8. Then:
  • (i)
    $0\leq\sup_{f}\widehat{\mathtt{D}}_{Z}(f)=\mathtt{D}_{\mathrm{KL}}\big{(}P_{\mathcal{M},B}\big{\|}\widetilde{Q}_{\mathcal{M}}\big{)}$.
  • (ii)
    For every $f$, $\widehat{\mathtt{D}}_{Z}(f)\leq\mathtt{D}_{\mathrm{KL}}\big{(}P_{\mathcal{M},B}\big{\|}\widetilde{Q}_{\mathcal{M}}\big{)}\equiv\mathtt{D}_{Z}$.
  • (iii)
    With probability at least $1-{\hat{\delta}}$ (over the independent validation splits in Definition 8), $0\leq\widehat{\mathtt{D}}_{\mathrm{LCE}}\leq\mathtt{D}_{Z}$.
Theorem 4 (SWD-Based Correction).
Let $Z=\mathcal{M}(X)+B$ and $\widetilde{Q}_{\mathcal{M}}=\mathcal{N}(\mu_{Z},\Sigma_{Z})$, and let $\lambda_{\max}(\Sigma_{Z})$ be the largest eigenvalue of $\Sigma_{Z}$. Assume $P_{\mathcal{M},B}\ll\widetilde{Q}_{\mathcal{M}}$ and $\Sigma_{Z}\succ 0$. Define
$\widehat{\mathtt{D}}_{Z}\equiv\frac{1}{2\lambda_{\max}(\Sigma_{Z})}\mathrm{SW}_{2}^{2}\big{(}P_{\mathcal{M},B},\widetilde{Q}_{\mathcal{M}}\big{)}$.
Then $0\leq\widehat{\mathtt{D}}_{Z}\leq\mathtt{D}_{Z}$.
Theorems 3 and 4 lead to Corollary 1.
Corollary 1.
Let $\mathcal{M}:\mathcal{X}\mapsto\mathbb{R}^{d}$ be an arbitrary deterministic mechanism and $B\sim\mathcal{N}(0,\Sigma_{B})$ such that $\mathtt{LogDet}(\mathcal{M}(X),B)=\beta$. Under the assumptions of Theorems 3 and 4, the perturbed mechanism $Z=\mathcal{M}(X)+B$ is PAC private with
$\mathtt{MI}(X;Z)\leq\beta-\widehat{\mathtt{D}}_{Z}<\beta,$ (8)
where $\widehat{\mathtt{D}}_{Z}>0$ is obtained by Theorem 3 ($\widehat{\mathtt{D}}_{Z}(f)$) or Theorem 4. In addition, if $\widehat{\mathtt{D}}_{Z}=\widehat{\mathtt{D}}_{\mathrm{LCE}}$, then (8) holds with probability at least $1-\hat{\delta}$.
Corollary 1 shows that accounting for non-Gaussianity through the correction term $\widehat{\mathtt{D}}_{Z}>0$ yields $\mathtt{MI}(X;Z)\leq\beta-\widehat{\mathtt{D}}_{Z}<\beta$, where the correction is obtained via DV-based correct or sliced Wasserstein correct. In practice, $\widehat{\mathtt{D}}_{Z}$ estimates the Gaussianity gap $\mathtt{Gap}_{d}$, capturing the saved privacy budget, which is particularly valuable for budget savings in mechanism composition. However, this budget-saving approach is post-hoc after Auto-PAC privatization. Appendix E provides additional discussions and interpretations. Next, we propose a new privacy framework enabling automatic optimal privacy budget implementation.
@@ 4 Residual-PAC Privacy
Recall that PAC Advantage Privacy (Definition 3) quantifies the amount of privacy leaked by $\mathcal{M}(X)$ in terms of the posterior advantage $\Delta_{f}^{\delta}$ encountered by the adversary. Complementing this perspective, we introduce the notion of posterior disadvantage encountered by the adversary, which captures the amount of residual privacy protection that persists after leakage by $\mathcal{M}(X)$.
To formalize this residual protection, we first define the intrinsic privacy of a data distribution $\mathcal{D}$ relative to a fixed reference distribution $\mathcal{R}$ on $\mathcal{X}$ such that (i) $\mathrm{supp}(\mathcal{D})\subseteq\mathrm{supp}(\mathcal{R})$ and (ii) the f-divergence $\mathtt{D}_{f}(\mathcal{D}\|\mathcal{R})$ is finite (when $\mathtt{D}_{f}$ is the KL-divergence, this means the entropy of $\mathcal{R}$ is finite; see Section 4.1 for the formal definition of Shannon/differential entropy). We then define
$\mathtt{IntP}_{f}(\mathcal{D})=-\,\mathtt{D}_{f}(\mathcal{D}\,\|\,\mathcal{R}),$
@@
where $\mathtt{D}_{f}(\mathcal{D}\|\mathcal{R})$ is the $f$-divergence between $\mathcal{D}$ and $\mathcal{R}$, quantifying how much $\mathcal{D}$ deviates from the reference $\mathcal{R}$. Intuitively, $\mathtt{D}_{f}(\mathcal{D}\|\mathcal{R})$ rewards distributions that remain close to the "random guess" using $\mathcal{R}$, and by construction $\mathtt{IntP}_{f}(\mathcal{D})\leq 0$, attaining zero exactly when $\mathcal{D}=\mathcal{R}$.
Examples of $\mathcal{R}$. When $\mathcal{X}$ is bounded, $\mathcal{R}$ can be the uniform law $\mathcal{U}$ on $\mathcal{X}$. However, on an unbounded $\mathcal{X}$, the uniform reference $\mathcal{R}=\mathcal{U}$ has infinite volume $\int_{\mathcal{X}}dx=\infty$, potentially making $\mathtt{IntP}_{f}(\mathcal{D})$ vacuous or undefined. To avoid this, we instead require $\mathcal{R}$ to satisfy $\mathtt{D}_{f}(\mathcal{D}\|\mathcal{R})<\infty$. For example, one can choose $\mathcal{R}$ by: (i) truncated uniform on a large but bounded region containing $\mathrm{supp}(\mathcal{D})$, (ii) maximum-entropy Gaussian matching known moments of $\mathcal{D}$, or (iii) smooth pullback of uniform on $(0,1)^{d}$ via bijection (e.g., component-wise sigmoid). Under any of these constructions, $\mathcal{R}$ retains the "random-guess" semantics yet has finite $\mathtt{D}_{f}(\mathcal{D}\,\|\,\mathcal{R})$, ensuring $\mathtt{IntP}_{f}(\mathcal{D})$ remains meaningful even on unbounded $\mathcal{X}$. Please see Appendix D for a detailed discussion.
Definition 9 ($(\mathtt{R}_{f}^{\delta},\rho,\mathcal{D})$ Residual-PAC (R-PAC) Privacy).
A mechanism $\mathcal{M}$ is said to be $(\mathtt{R}_{f}^{\delta},\rho,\mathcal{D})$ Residual-PAC (R-PAC) private if it is $(\delta,\rho,\mathcal{D})$ PAC private and
$\mathtt{R}_{f}^{\delta}\equiv\mathtt{IntP}_{f}(\mathcal{D})-\mathtt{D}_{f}(\mathbf{1}_{\delta}\|\mathbf{1}_{\delta^{\rho}_{o}}),$
@@
is the posterior disadvantage, where $\mathbf{1}_{\delta}$ and $\mathbf{1}_{\delta^{\rho}_{o}}$ are indicator distributions representing the adversary’s inference success before and after observing the mechanism’s output, respectively.
@@
The posterior disadvantage $\mathtt{R}_{f}^{\delta}$ captures the residual privacy guarantee, which is the portion of intrinsic privacy (w.r.t. a reference $\mathcal{R}$) that remains uncompromised after the privacy loss $\Delta_{f}^{\delta}=\mathtt{D}_{f}(\mathbf{1}_{\delta}\|\mathbf{1}_{\delta^{\rho}_{o}})$ (Definition 3). Then, the total intrinsic privacy is precisely decomposed as
$\mathtt{IntP}_{f}(\mathcal{D})=\mathtt{R}_{f}^{\delta}+\Delta_{f}^{\delta}.$ (9)
@@
This relationship provides a complete and interpretable quantification of privacy risk, distinguishing between the privacy that is lost and that which endures after information disclosure via $\mathcal{M}(X)$. Analogous to PAC Privacy, membership inference attacks (MIA) and R-PAC Membership Privacy can be instantiated from R-PAC Privacy. See Appendix B for detailed constructions.